A circular coil of wire has a diameter of 20.0 cm and contains 10 loops. The current in each loop is 3.00 amps, and the coil is placed in a 2.00-T magnetic field. What are the minimum and maximum torques exerted on the coil by the magnetic field?

Please show me how you got the answer
Answer #1

Tmin = NIAB sin theta

for min thta = 0 deg

so T = 0

B>

torque T = NiAB sin 90

where N = No. of loops = 10

A = area    pir ^2 here r= d/2 = 20/2 = 10 cm = 0.1 m

i = current i = 3A

B = magnetic field = 2 T

T = 10 * 3 * 0.1*0.1* 2*1

Tmax = 1.884 Nm
Answer #2

Part A)

The minimum is easy, that would be zero

The max is from T = NBIA

Tmax = (10)(2)(3)(pi)(.1)2

Tmax = 1.88 Nm